OpenStudy (anonymous):

Which is the correct classification for this system? x - y = 1 2x + y = 4 A. Inconsistent B. Consistent independent C. Coincident

jimthompson5910 (jim_thompson5910):

add the two equations x-y = 1 2x+y = 4 --------- 3x+0y = 5

jimthompson5910 (jim_thompson5910):

3x+0y = 5 turns into 3x = 5 which means x = ???

OpenStudy (anonymous):

1.6

jimthompson5910 (jim_thompson5910):

then we can use this x value to find the corresponding y value

jimthompson5910 (jim_thompson5910):

x-y = 1 1.6-y = 1 y = ??

OpenStudy (anonymous):

0.6???

OpenStudy (anonymous):

so B?

jimthompson5910 (jim_thompson5910):

yep we have a single solution so the system is both independent and consistent

OpenStudy (anonymous):

thanks(:

jimthompson5910 (jim_thompson5910):

yw